OpenStudy (anonymous):

If x = 1 is a zero of the polynomial function f(x) = 3x3 − 8x2 + 3x + 2, which of the following is another zero of f(x)? MEDAL AND FAN PLEASE

OpenStudy (anonymous):

@welshfella @WazapxD @Michele_Laino

OpenStudy (anonymous):

wouldnt be a x behind 2

OpenStudy (anonymous):

4x^3-8x^2+3x+2

OpenStudy (anonymous):

ima work this out brb

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whats the answer choices

OpenStudy (anonymous):

(x-2) (3x+1) (x-1) are the factors of the polynomial function

OpenStudy (anonymous):

sets of zeroes are x=2 x= -1/3 x=1

OpenStudy (anonymous):

The answer choices are 2, -2, -1, 5

OpenStudy (anonymous):

it is 2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

you need to factor the trinomial function to find the zeroes

OpenStudy (anonymous):

but wouldn't it be -1,2, and 1/3

OpenStudy (anonymous):

yes that is correct they're 3 zeroes in a cube root unless one of the zeroes is repeated

OpenStudy (anonymous):

they are only looking for another one

OpenStudy (anonymous):

which is 2 or -1/3 since 1 is given

OpenStudy (anonymous):

ok. can i ask another?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Anyone?

OpenStudy (welshfella):

this is the Descartes rule of signs which i've only got a vague memory of . I'll see if I can find a website for you. Or maybe someone who remembers it will come along..

OpenStudy (anonymous):

Here is a link to a similar question

OpenStudy (anonymous):

But I still don't understand it :(

OpenStudy (welshfella):

http://www.purplemath.com/modules/drofsign.htm

OpenStudy (anonymous):

:) Thank you!!

OpenStudy (anonymous):

Which polynomial identity will prove that (3 + 6)2 = 81?

OpenStudy (anonymous):

power rule

OpenStudy (welshfella):

there are 3 changes of sign in this polymomial

OpenStudy (anonymous):

so would it be difference of squares?

OpenStudy (anonymous):

Square of a Binomial

OpenStudy (anonymous):

Which of the following is a polynomial with roots: −square root of 3, square root of 3, and −2?

OpenStudy (welshfella):

now you find f(-x) and count the sign changes again f(-x) = (-x)^6 −(- x)^5 − (-x)^4 + 4(-x)^3 − 12(-x)^2 + 12 = x^6 + x^5 - x^4 - 4x^3 -12x^2 + 12 - 2 changes of signs

OpenStudy (anonymous):

So for the positive would it be 4,2,0 sign changes?

OpenStudy (welshfella):

i missed one out for f(x) its 4 sign changes

OpenStudy (welshfella):

not sure - i'll have to re check

OpenStudy (anonymous):

Ok. So would the answer be positve:4,2,0 and negative 2,0 and complex:6,4,2,0

OpenStudy (welshfella):

yea 4,2 or 0 positive roots

OpenStudy (anonymous):

Which of the following is a polynomial with roots: −square root of 3, square root of 3, and −2?

OpenStudy (anonymous):

x3 − 2x2 − 3x + 6 x3 − 3x2 − 5x + 15 x3 + 2x2 − 3x − 6 x3 + 3x2 − 5x − 15

OpenStudy (anonymous):

is the answer B?

OpenStudy (welshfella):

yes i think you are right for the first one

OpenStudy (anonymous):

so is it B for the second question?

OpenStudy (welshfella):

roots sqrt3 - sqrt3 and -2 factors will be (x + 2)(x - sqrt3)(x + sqrt3) = (x +2)(x^2 - 3)

OpenStudy (welshfella):

expanding gives x^3 + 2x^2 - 3x - 6

OpenStudy (anonymous):

So it would be A?

OpenStudy (welshfella):

that is C

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

x3 + 2x2 − 3x − 6

OpenStudy (anonymous):

That one?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

I thought it was A. But i believe u

OpenStudy (anonymous):

What is the quotient (6x4 − 15x3 + 10x2 − 10x + 4) ÷ (3x2 + 2)?

OpenStudy (anonymous):

2x2 − 5x + 2 I got this as an answer

OpenStudy (anonymous):

yes?

OpenStudy (welshfella):

sorry I got called away lets see

OpenStudy (anonymous):

It's okay:)

OpenStudy (anonymous):

What is the quotient (6x4 − 15x3 + 10x2 − 10x + 4) ÷ (3x2 + 2)? Just in case u needed the question again

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

looks good

OpenStudy (anonymous):

Was I right?

OpenStudy (welshfella):

lol - i'm really slow with this one

OpenStudy (anonymous):

it's fine:)

OpenStudy (welshfella):

Yes you are right

OpenStudy (anonymous):

ok Using synthetic division, what is the quotient (2x3 − 2x − 12) ÷ (x − 2)?

OpenStudy (anonymous):

I got this as an answer: 2x2 + 4x + 6

OpenStudy (welshfella):

correct

OpenStudy (anonymous):

ok If x = 1 is a zero of the polynomial function f(x) = 3x3 − 8x2 + 3x + 2, which of the following is another zero of f(x)?

OpenStudy (anonymous):

i got -1

OpenStudy (welshfella):

f(-1) = -3 - 8 - 3 +2 = -12 so its not -1 as the last term is + 2 it might be -2 or 2

OpenStudy (anonymous):

both -2, and 2 are answer choices :(

OpenStudy (welshfella):

try f(2)

OpenStudy (welshfella):

f(2) = 3(8)- 8(4) + 6 + 2 = 24 - 32 + 8 = 0

OpenStudy (welshfella):

so its 2

OpenStudy (anonymous):

What is the remainder when (3x4 + 2x3 − x2 + 2x − 24) ÷ (x + 2)?

OpenStudy (anonymous):

Thank-you:) For this one I got 0 as a answer

OpenStudy (welshfella):

you should really be submitting them in different posts .. but ill do this last one

OpenStudy (anonymous):

Oh okay. I'm sorry :)

OpenStudy (welshfella):

f(-2) will give you the remainder

OpenStudy (anonymous):

so it's not 0?

OpenStudy (welshfella):

3(-2)^4 + 2(-2)^3 - (-2)^2 - 4 - 24 = 48 - 16 - 4 - 4 - 24 = 48 - 48 = 0

OpenStudy (welshfella):

correct